# Expert Study on Interpretable Machine Learning Models with Missing Data

### v1

@@ Conclusion
We would like to express our great gratitude to the healthcare professionals who participated in the experiment. We appreciate that they voluntarily devoted their time to complete the experiment. This paper and the research behind it would not have been possible without the significant feedback they delivered.
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A
A.1 Survey setup
@@
Survey instrument
The survey was developed in English with input from trauma physicians and ML researchers and then translated into French by native speakers for accuracy and cultural relevance. The 41 questions included demographics, single/multiple-choice items with comments, ranking tasks, open-ended questions, and statement blocks on a 5-point Likert scale (Taherdoost, 2019). Likert scales gauged attitudes, choice questions collected structured data, ranking tasks prioritized preferences, and open-ended questions offered deeper insights. Appendix A.3 contains the questionnaire in both languages.
@@
Figure 2: Experimental setup. Physicians are shown a data entry of a patient with 9 features in an interpretable machine-learning model along with the questions. After the data is gathered qualitative and quantitative methods are used to analyze the results.
A.2 Use case
We aim to understand the challenges of diagnosing hemorrhagic shock in trauma patients, a leading cause of early, preventable death in cases of severe injury (Kauvar and Wade, 2005; Dutton et al., 2010). Timely intervention is crucial, as delays can significantly worsen prognosis. For our survey, hemorrhagic shock is defined as a composite outcome that includes any patient who meets one or more of the following criteria: 1) administration of at least one packed red blood cell (RBC) in the resuscitation room, 2) transfusion of four or more RBC units within six hours of admission, 3) the need for a hemorrhage control procedure (such as interventional radiology or surgery) within six hours of admission, or 4) death from hemorrhagic shock within 24 hours (Hamada et al., 2018). The data in the use case include both continuous and categorical features, covering clinical variables like heart rate and non-clinical factors such as prior clinician decisions, including ordering specific lab tests.
A.3 Survey questions
@@
A.4 Additional results
This section presents further findings, including figures on the attitudes of different cohorts and the characteristics of identified clusters within these cohorts. Additionally, we provide detailed rankings of participants’ preferences for various imputation methods across different IML techniques.
Cohort analysis
We conducted a Factor Analysis for Mixed Data (FAMD) Josse and Husson (2012), revealing four clusters among the survey responses. Missing values e.g. unanswered questions, were imputed using Principal Component Analysis (PCA) (Josse and Husson, 2012) for continuous variables and the most frequent category for categorical data. The primary goal FAMD Josse and Husson (2012) and the clustering is descriptive—to provide a multivariate view of clinician groups rather than relying on just one or two variables. We used FAMD Pagès (2014); Escofier and Pages (1994), a low-rank approximation method specifically suited for mixed data (both continuous and categorical), serving as the counterpart to PCA Greenacre et al. (2022) for such datasets. FAMD is particularly useful for analyzing questionnaire results, as it balances the influence of both quantitative and categorical variables while computing distances between observations (in this case, clinicians). The first principal component of FAMD is a quantitative variable most related to all others, measured by squared correlation for continuous variables and $R^{2}$ for categorical variables (similar to how PCA works for quantitative data , with $R^{2}$ analogous to an ANOVA model for categorical data).
We identified four clusters (Figure in Appendix LABEL:fig:cohort_cluster) based on clinicians’ attitudes towards AI/ML, their familiarity and trust in IML, and their preferences for handling missing values, regardless of demographic factors. Cluster 1 includes clinicians (16% of 55, 9 people) with low understanding and use of risk scores, little familiarity with decision trees, and dissatisfaction with all three models when dealing with missing values. Despite this, they are comfortable using black-box models like MICE but prefer implicit imputation and surrogate splits. Cluster 2 consists of clinicians (22%, 12 people) who show high levels of fear of AI, blind trust in its capabilities, and concerns about job loss, alongside perceiving current AI solutions as low-quality. They are familiar with decision trees and favor zero imputation or surrogate splits. Cluster 3 has substantial missing demographic data and may need to be excluded from the analysis (36%, 20 people). Cluster 4 comprises clinicians (25%, 14 people) who use AI daily, are highly familiar with all three IML models, and feel confident explaining these models with or without missing values. They favor implicit imputation over black-box models and surrogate splits.
@@ A.4 Additional results
app:tab:stats_participants Characteristic Total (N=55) Complete answers (N=20) Age in years, mean (SE) 41,18 (1,13) 39,35 (1,55) Female, n (%) 22 (40%) 7 (5%) Male, n (%) 33 (60%) 13 (65%) Highest received educational degree, n (%)    University bachelor/master/postgraduate education 9 (16,36%) 3 (15%)    Medical doctoral degree 37 (67,27) 12 (60%)    Associate Professor/Professor 9 (16,36%) 5 (25%) Current level of occupation, n (%)    Clinical doctor 46 (83,64%) 19 (95%)    Intern/Trainee/Non-fully graduated clinician 3 (5,45%) 0    Other 6 (10,91%) 1 (5%) Medical specialty, n (%)    Anesthesiology /// Intensive Care 46 (83,64%) 16 (80%)    Emergency medicine 5 (9,09%) 2 (10%)    Other 4 (7,27%) 2 (10%)
Table 3: Descriptive statistics of survey participants. Categorical variables have been simplified for brevity: The ’Other’ category for current occupation included roles such as nurse, clinical studies technician, hospital clinical trials coordinator, clinical manager assistant, and retired professional. For medical specialization, it included fields like rehabilitation, pediatrics, surgery, anesthesiology, bioinformatics, pediatric orthopedic surgery, and clinical research associate.
[…]
tab:Imputation_ranking
Table 4: The table presents the results of clinicians’ rankings of implicit and explicit imputation methods across different IML techniques. Clinicians were asked to evaluate and rank three IML methods, focusing on their preferences between implicit imputation, which considers other feature values for handling missing data and allows for expert adjustments, and explicit imputation, which uses methods like zero or mean imputation. The table also includes rankings for more complex and less transparent algorithms, such as Multivariate Imputation by Chained Equations (MICE) (Van Buuren and Groothuis-Oudshoorn, 2011), used to fill in missing data.
Rank which method—decision tree or risk score—you think implicit imputation is more suitable for?
Rank Risk Scores Decision Tree Sum
Rank 1 9 8 17
Rank 2 8 9 17
For which interpretable model would you most likely choose the black box imputation?
Rank Risk Scores Decision Tree Coefficients in Linear Models
Rank 1 4 4 7
Rank 2 3 5 7
Rank 3 6 4 5
Sum 13 13 19

### v2

@@ 3.1 Finding 1: Clustering based on AI/ML attitudes and missing values practices
tab:stats_participants Characteristic Total (N=55) Complete answers (N=20) Age in years, mean (SE) 41.18 (1.13) 39.35 (1.55) Female, n (%) 22 (40%) 7 (5%) Male, n (%) 33 (60%) 13 (65%) Highest received educational degree, n (%)    University bachelor/master/postgraduate education 9 (16,36%) 3 (15%)    Medical doctoral degree 37 (67,27%) 12 (60%)    Associate Professor/Professor 9 (16,36%) 5 (25%) Current level of occupation, n (%)    Clinical doctor 46 (83,64%) 19 (95%)    Intern/Trainee/Non-fully graduated clinician 3 (5,45%) 0    Other 6 (10,91%) 1 (5%) Medical specialty, n (%)    Anesthesiology /// Intensive Care 46 (83.64%) 16 (80%)    Emergency medicine 5 (9.09%) 2 (10%)    Other 4 (7,27%) 2 (10%)
Table 1: Descriptive statistics of survey participants. Categorical variables have been simplified for brevity: The ’Other’ category for current occupation included roles such as nurses, clinical studies technicians, hospital clinical trials coordinators, clinical manager assistants, and retired professionals. For medical specialization, it included fields like rehabilitation, pediatrics, surgery, anesthesiology, bioinformatics, pediatric orthopedic surgery, and clinical research associate.
We identified four distinct clusters among survey participants based on the FAMD results (see Section 2). This provides a structured understanding of how clinicians’ attitudes toward AI/ML and their handling of missing values vary across different response groups (Figure LABEL:fig:attitude_for_AI_ML and  LABEL:fig:cohort_cluster in the Appendix). Cluster 1 includes clinicians (16% of 55, 9 participants) with low understanding and use of risk scores, little familiarity with decision trees, and dissatisfaction with all three models when dealing with missing values. Despite this, they are comfortable using black-box models like MICE but prefer implicit imputation and surrogate splits. Cluster 2 consists of clinicians (22%, 12 people) who show high levels of fear of AI, blind trust in its capabilities, and concerns about job loss, alongside perceiving current AI solutions as low-quality. They are familiar with decision trees and favor zero imputation or surrogate splits. Cluster 3 (36%, 20 people) has significant missing data and may be excluded. Cluster 4 comprises clinicians (25%, 14 people) who use AI daily, are highly familiar with all three IML models, and feel confident explaining these models with or without missing values. They favor implicit imputation over black-box models and surrogate splits. These clusters highlight the diverse perspectives among clinicians, emphasizing the challenge of developing approaches that accommodate varying familiarity, trust, and confidence in handling missing data.
Current practices with missing values.
Across all clusters, we identified various strategies to manage missing values in currently used clinical scores, e.g. RedFlag (Dutton et al., 2010), primarily relying on clinical intuition or other available clinical information. Many mentioned using plausible values based on clinical context or the patient’s history to estimate missing values. Some prefer worst-case imputation to avoid underestimating risk, though one noted it may lead to overtraining and resource inefficiency. Others choose to partially calculate or omit scores if key data is missing, feeling that imputation could ‘skew the score concept’. One clinician also suggested using alternatives, like thromboelastography when prothrombin time is unavailable.
3.2 Finding 2: Clinician preferences and interpretations from a use case with IML approaches and missing values
Next, we analyze clinicians’ responses from the use case evaluation, focusing on their approach to handling missing patient data in IML. We summarize their interpretation patterns and preferences for imputation or native methods.
Confidence does not always align with accuracy when handling missing values.
Out of 20 responses, 9 clinicians rated themselves as highly confident on a Likert scale in explaining risk scores, even with missing values, while 10 expressed the same level of confidence for decision trees. For logistic regression, however, confidence was lower and accompanied by uncertainty; only 6 participants were correct, with 4 out of 6 correct among those expressing confidence in its usefulness with missing values.
Low trust in zero and mean imputation.
When clinicians received the mean hemoglobin value for the dataset, almost all 23 respondents reported that this additional information neither influenced their decisions nor supported their decision-making in all three IMLs. The main reasons included the irrelevance to individual cases: “I don’t take it into account. Each patient is unique, and I don’t see the relevance of giving him an average value.” When presented with a prediction for a patient in which the missing value was imputed as zero, most clinicians indicated that they would ignore the feature in the risk score. Similarly, a machine learning model would also ignore the feature, effectively diminishing its influence on the prediction. While the representation of risk scores was still considered helpful by most, some clinicians preferred more accurate imputations for the linear model, with one remarking: “If imputation is used during testing, it should be clearly indicated, ideally with the associated uncertainty. This transparency allows for comparison between predictions with and without the imputed value.”
Surrogate splits can be viable.
For decision trees, using surrogate splits (the direction to take when the splitting feature is missing) was explored, but this method was less familiar to clinicians and caused some confusion. However, many clinicians noted that its usefulness depends on clearly indicating where the substitution occurs: “If the substitution is clearly indicated and can be taken into account by the clinician, yes, it could be a decision aid.” Despite this, some preferred other methods, one of them stating: “Decision trees are easy to understand, but unstable and too sensitive to values. That’s why we prefer Random Forest algorithms. It’s difficult to rely on a decision tree, especially if there are missing values that have been imputed Rather than imputing […], there are algorithms that estimate missing values in relation to other parameters.”
High trust in risk scores and decision trees despite missing values.
Lastly, the survey results show varying levels of trust among the 23 clinicians in using the three IML models for diagnosing hemorrhagic shock. Risk Scores were the most trusted, with 15 clinicians indicating they ’Almost Trust’ the model. However, concerns about over-reliance on numbers rather than clinical judgment were noted, as one clinician remarked, “I’m not dealing with numbers, but with a patient! I need a history, background treatment, a clinical history […]” Linear models faced the most skepticism, with 15 clinicians ’Not Sure’ or ’Rather not’ trusting them, due to unfamiliarity and doubts about handling missing values. Decision trees had mixed feedback—13 clinicians ’Fully or Almost Trust’ them for their clarity, but some raised concerns about potential limitations.
Comparison of imputation methods across IMLs.
[…]
fig:imputation_IML Refer to caption
Figure 3: Clinician preferences for imputation methods across different IML models. We normalize by dividing the number who chose a combination by the total, as the total votes for a model can vary.
Clinically reasoned preferences for missing values at test time.
When choosing between implicit and explicit imputation methods, participants strongly preferred implicit imputation. This preference stems from the need for transparency, ease of understanding, and alignment with clinical reasoning, crucial in medical decision-making. One participant emphasized these factors, stating, “I prefer implicit imputation, as I’d find it hard to trust a black-box algorithm. Ease of understanding and communication are important, as well as from a ’medico-legal’ point of view, to ultimately make a decision that involves my responsibility and the patient’s life! […] Even if the computer makes the decision, the responsibility […] lies with me, and I must be able to understand and justify it.” This highlights the importance of comprehensibility in high-stakes settings like healthcare. Another respondent noted, “I prefer implicit imputation because it seems simpler to reason with the values we have than to rely on a complex algorithm we don’t understand.” This reflects broader concerns among clinicians about the trustworthiness and transparency of black-box imputation. The practical aspects of clinical work were also highlighted by a participant who remarked, “Implicit imputation is more comprehensible and easier for clinicians to understand. Nevertheless, we need to see which technique is the best in terms of prediction, and I can’t say what I prefer without knowing the relevance and PPV/VPn of each technique.” This illustrates the balance clinicians must strike between understanding a method and ensuring its effectiveness. Finally, the intuitive nature of implicit imputation was underscored by another participant, who stated, “I prefer implicit imputation because, in my opinion, it’s the closest thing to clinical reasoning: […] it’s the most intuitive and perhaps the safest approach, rather than an arbitrary rule.” Some responses suggest sophisticated methods like MICE improve predictive performance by better preserving variable relationships than simpler methods like mean or zero imputation.
3.3 Finding 3: Requirements for designing IML models for clinical integration
Most respondents were open to using IML for diagnosis with missing values, though some were uncertain. A key concern was the need for clear guidelines on handling missing values. Clinicians emphasized the importance of structured explanations, with one suggesting that models should present multiple imputed values for intuitive selection, while another stressed the need for “a clear, simple, and rational explanation of how this missing data is treated, and its consequences on the test result.” Without such transparency, trust in IML recommendations remains limited. Understanding and communicating uncertainty also emerged as a critical factor. Clinicians highlighted the challenge of making decisions with missing values and emphasized the importance of incorporating decision thresholds. One respondent noted that they could make a decision despite missing values but would “weight my decision with the patient’s clinical data,” underscoring the need for models to reflect confidence levels in their outputs. Beyond uncertainty, model reliability remains a fundamental requirement. Clinicians stressed the importance of ensuring that both models and imputation methods produce stable and trustworthy results across different clinical scenarios. Without this assurance, the adoption of IML in practice would be challenging. Equally important is the need for interpretability and accessibility, particularly for non-technical users. Some clinicians highlighted the value of external validation and visual aids to clarify how missing values impact model outputs, making results easier to understand and act upon. Finally, clinical intuition must remain central to decision-making. Clinicians emphasized that IML should enhance, rather than replace, human expertise. One respondent stressed that “the human element and intuition based on experience and the CLINICAL EXAMINATION OF THE PATIENT!!!!” should continue to guide clinical decisions, highlighting the need for models that complement expert judgment rather than override it.
4 Related work
Our study aligns with previous findings that AI positively influences healthcare decision-making. van der Meijden et al. (2023) show in their study, that 97% of respondents were familiar with AI, and 86% believed it could aid clinical work. However, concerns remain about confirmation bias, over-reliance on models, and the effort needed for tool interaction (Wysocki et al., 2023). Our results suggest readiness for IML models, but successful implementation requires clinician participation in design to address needs like understanding missing data mechanisms and uncertainties (Lage et al., 2019a, b). While Janssen et al. (2010) argues that simple methods for handling missing data can lead to misleading results, others advocate for multiple imputation techniques as they enhance validity in clinical research (Austin et al., 2021). In contrast, our expert evaluation suggests implicit imputation, based on clinical intuition, is generally preferred at test time, while black-box models are the least favored due to interpretability issues. Notably,  Read (2024) highlights that senior critical care physicians have traditionally relied on autonomy and clinical intuition. While this approach enables adaptability in handling complex and unpredictable scenarios, it also contributes to unwarranted variations in care delivery and potential harm. Following Tonekaboni et al. (2019), who emphasized evaluating clinical explainability in high-stakes settings, IML models must undergo rigorous validation to assess performance, interpretability, and legal implications (Wiens et al., 2019). Clinicians can build trust in IML models by understanding how they work and through positive experiences where outputs align with their domain knowledge (Kelly et al., 2019). Randomized controlled trials can validate models but are costly, time-consuming, and may lack generalizability (Shortliffe and Sepúlveda, 2018). Our findings suggest that successful deployment requires robust and purpose-built models, a clear presentation of uncertainties (Doshi-Velez and Kim, 2017), and training clinicians on the impact of missing values to integrate these tools into workflows (Topol, 2019; Maddox et al., 2019).
5 Discussion and conclusion
Our survey evaluates how clinicians interact with predictive models, particularly in handling missing values. Our findings reveal a disconnect between missing data handling in ML and retrospective research, such as MICE imputation, and clinical decision-making, where clinicians rely more on observed data and medical intuition. While imputation methods like zero, mean, or black-box approaches may aid model training, they are less favored at test time. Worst-case imputation, though clinically cautious, can introduce statistical bias and misrepresent patient risk levels.
To better align with clinical needs, future methods should either reduce reliance on imputation or adopt more transparent strategies. Native handling of missing values, as seen in tree-based models, can, for instance, be implemented using ”Missingness Incorporated in Attributes” (MIA), which treats missing values as a distinct category Twala et al. (2008); Kapelner and Bleich (2015); Josse et al. (2024). This approach is similar to adding missingness indicators in linear models, although increasing the number of features can reduce interpretability. Stempfle and Johansson (2024), a generalized linear rule model minimizing reliance on missing values, avoids the need for imputation during training or testing, offering an intriguing direction for future research.
The diversity of clinician attitudes toward AI highlights the challenge of aligning models with clinical intuition and user preferences. Clinicians in Clusters 1 and 2, unfamiliar with IML, fearful of AI, or reliant on black-box models, would benefit from training on interpretable models, handling missing data, and AI’s supportive role in practice. In contrast, Cluster 4, comprising highly engaged AI users, is well-suited for co-designing IMLs and refining missing value strategies. Capturing clinician intuition (Read, 2024) by integrating preferred practices in a way similar to computer programming holds great potential, particularly in complex or evidence-limited scenarios. Without clinician involvement, the adoption of these tools is unlikely, underscoring the need for IML methods that enhance clinician interaction with predictive models, particularly for handling missing values.
This study lays the foundation for understanding the challenges in deploying clinical prediction models with missing values. However, it is subject to self-selection bias due to clinicians’ demanding workloads and limited availability. Future work could explore how these findings apply to other learning paradigms, like reinforcement learning, and different clinical settings, such as outpatient care,emphasising on the interaction between IMLs and clinicians. Expanding the study beyond the size of clinicians in France, we could also extend the study to include other healthcare systems for broader insights.
Appendix A Survey details
Survey setup
@@ Appendix A Survey details
Figure 4: Experimental setup. Clinicians are shown a data entry of a patient with 5 features whereas one feature is missing in an interpretable machine-learning model along with the questions. After the answers are gathered qualitative and quantitative methods are used to analyze the results.
Use Case of Hemorrhagic Shock
Diagnosing hemorrhagic shock is often complex, as there is no universally accepted minimal set of criteria. Research, including a study by the Traumabase teams (James et al., 2022), has highlighted the coexistence of multiple definitions of shock. In clinical practice, the diagnosis is based on a combination of factors indicating decreased cardiac output and hemorrhage. These include patient history (e.g., potential sources of bleeding), clinical examination (e.g., hypotension, tachycardia, visible bleeding), and paraclinical tests (e.g., active bleeding on CT scan or ultrasound findings). The survey use case was designed with input from trauma specialists, and we can clarify this further in the paper.
Missing values in FAMD
The primary goal of the Factor Analysis for Mixed Data (FAMD) (Josse and Husson, 2012) analysis and clustering is descriptive, allowing us to characterize clinicians from a multivariate perspective, rather than relying on just one or two variables to uncover patterns and group-specific insights. FAMD is the counterpart of PCA for mixed data, both continuous and categorical and is well-suited for describing results from questionnaires. Excluding incomplete responses could introduce bias if the missing data affects a specific subpopulation. Therefore, we compute the principal components using an EM algorithm, which maximizes the observed likelihood. In low-rank approximation methods like PCA or FAMD, estimation is equivalent to missing values and imputation. This means that minimizing the weighted least squares criterion or maximizing the observed likelihood yields the same results. The algorithms in missMDA (Josse and Husson, 2016) use an iterative SVD (Hastie et al., 2015) process: impute, compute SVD, update the imputation, and repeat. Thus, the imputation is ”neutral” because the SVD on the imputed dataset produces the same results as on the dataset with missing values.
Appendix B Survey questions
@@
Appendix C Additional results
This section presents further findings, including figures on the attitudes of different cohorts and the characteristics of identified clusters within these cohorts.
@@ Appendix C Additional results
fig:attitude_for_AI_ML Refer to caption
Figure 5: Cohort’s attitude on AI/ML. We show the mode and standard deviation for eight statements exploring perspectives on AI/ML tools in healthcare, covering familiarity with these technologies, beliefs regarding their potential to replace physicians, expectations about their added value and support, and whether these tools adequately represent physicians’ work to be useful. Missing values were not yet discussed.